Hi-Tek Labs performs steroid testing services for colleges and universities. Because the company works only with educational institutions, the price of each test is strictly regulated. Therefore, the costs incurred must be carefully monitored and controlled. Shown below are the standard costs for a typical test:

Direct materials (1 Petri dish at $1.80 per dish)

$ 1.80

Direct labour (0.5 hours at $20.50 per hour)

10.25

Variable overhead (0.5 hours at $8 per hour)

4.00

Fixed overhead (0.5 hours at $5 per hour)

  2.50

Total standard cost per test

$18.55

The lab does not maintain an inventory of Petri dishes. Therefore, the dishes purchased each month are used that month. Actual activity for the month of May 2020, when 2,500 tests were conducted, resulted in the following:

Direct materials (2,530 dishes)

$ 5,060

Direct labour (1,240 hours)

26,040

Variable overhead

10,100

Fixed overhead

5,700

Monthly budgeted fixed overhead is $6,000. Revenues for the month were $55,000, and selling and administrative expenses were $2,000.

Instructions

a.  

Calculate the price and quantity variances for direct materials and direct labour.

LQV $205 F

b.  

Calculate the total overhead variance.

c.  

Prepare an income statement for management.

d.  

Provide possible explanations for each unfavourable variance.
Answer #1

(a)   actual estimated

direct material: $5060/2530 dishes = $2 per dish $1.8 per dish (given)

i.e., $1.8per dish * 2530 dishes = $4554

price variance: $2 - $1.8 = $0.2 per dish

total  price variance: $5060 - $4554 = $506

total  price variance: $26040/1240 hours =$21 per hour $20.5 per hour (given)

  i.e.,$20.50 per hour * 1240 hrs =$25420

price variance: $21 - $20.5 = $0.5 per hour

total  price variance: $26040 - $25420 = $620

(b)

variable overhead: $10100/1240 hours= $8.14 $8 per hour (given)

$8 per hour * 1240 hours = $ 9920

fixed overhead: $5740/1240 hours = $4.59 $5 per hour

$5 per hour * 1240 hours= $6200

total overhead variance: $8.14 + $ 4.59 = $12.73 per hour    $8 + $5 = $13  per hour

i.e., $12.73 - $13 = (0.27) per hour under rated variance